Hi I need help with this linear algebra question. Thank you!

Let 0 A 3 0 9 3 -5 0 10 -2 Find eigenvalues of A and determine a basis for each eigenspace.
Answer #1

are Eigenvalues of AXI 0 1 3 اے a 3 lo = O -5 -2 3-) q 3-X = 0 10 -2-). -5 Q = (3-1) (3-1) 4-2-4) X=3, 3, -2 Hence are -2,3,3(0 o., 5). (o,- 29, 3) (о, -2, ) та Hence eigenspace span { lo, -2,1) eigenspace corresponding to X=3 (e-(5)x] [ () - S) - о